A 24-N horizontal force is applied to a 40-N block initially at rest on a rough horizontal

surface. If the coefficients of friction are µs = 0 .5 and µk = 0 .4, find the magnitude of the

frictional force on the block


1
Expert's answer
2021-10-21T16:34:16-0400

Frictional force is

"Ff = µk \times mg"

"Ff = µk \times (F) \\

Ff =0.4 \times (40) \\

Ff =16 m"